OpenStudy (anonymous):

Simplify: √(a^2 +2a +1)

OpenStudy (anonymous):

\[\sqrt{(a+1)^2}\]

OpenStudy (anonymous):

then that would just be a+1 ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Do you know how he got (a + 1) out of the initial equation though?

OpenStudy (anonymous):

yes... a^2 +2a +1 when factored is (a+1)^2

OpenStudy (anonymous):

Alright, just making sure :)

OpenStudy (anonymous):

thanks :)

OpenStudy (nikvist):

\[\sqrt{a^2+2a+1}=\sqrt{(a+1)^2}=|a+1|\]